Find the probability that the sum is as stated when a pair of dice is rolled.odd and less than 5
1
Expert's answer
2020-05-21T12:47:24-0400

Each die has six possible equiprobable states "x= 1,2,3\dots 6" and "y = 1,2,3\dots 6"

Since the dies are independent, "6*6 = 36" , unique states "(x,y)" are possible.

The possible states when the sum is odd and less than 5 are:

"(x,y) = \{(1,2),(2,1)\}" .

Only 2 states are possible.

Thus, the probability that the sum is odd and less than 5 is:

"\frac{\textrm{number of states when the sum is odd and less than 5} }{\textrm{number of all possible states}} = \frac{2}{36} = \frac{1}{18}" .